OpenStudy (anonymous):

I'm trying to figure out the p-series to compare to the infinite series 1/n + ln n. Is it 1/n or 1/n^2?

OpenStudy (anonymous):

\[\frac{1}{n}>\frac{1}{n+\ln(n)}>\frac{1}{n^2}\]

OpenStudy (anonymous):

so if I was doing a limit comparison test I would use the infnite series \[1/n^2\]?

OpenStudy (anonymous):

i don't think you can use it here

OpenStudy (anonymous):

ah ok thanks for the help